We want to estimate the mean weekly earnings of students at a particular college with 95% confidence. How many students must be randomly selected so that the sample mean is within $1 of the population mean? Population standard deviation is known to be $10.
Answer #1

Solution :

Given that,

standard deviation =  \sigma =10

Margin of error = E = 1

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

Z\alpha/2 = Z0.025 = 1.96

sample size = n = [Z\alpha/2* \sigma / E] 2

n = ( 1.96* 10 /1 )2

n =384.16

Sample size = n =384